# Vector Dark Matter with Higgs Portal in Type II Seesaw framework

### v1

@@ Abstract
We study the phenomenology of a vector dark matter (VDM) in a $U(1)_{X}$ gauged extension of the Standard Model (SM) which is connected to the type II seesaw framework via the Higgs portal. When this $U(1)_{X}$ symmetry is spontaneously broken by the vacuum expectation value (VEV) of a complex scalar singlet, the gauge boson $Z^{\prime}$ becomes massive. The stability of the dark matter (DM) is ensured by the introduction of an exact charge conjugation symmetry. On the other hand, the $SU(2)_{L}$ triplet scalar generates light neutrino masses through the type II seesaw mechanism. We have studied the phenomenology of the usual WIMP DM considering all possible theoretical and experimental constraints that are applicable. Due to the presence of triplet scalar, our scenario can accommodate the observed $2\sigma$ deviation in $h\to Z\gamma$ decay. We have also briefly discussed the possibility of non-thermal production of DM from the decay of the same complex scalar that is responsible for the breaking of this $U(1)_{X}$ symmetry.
@@ I Introduction
The existence of a non-luminous, non-baryonic form of matter which provides roughly $26\%$ energy density of the universe, popularly known as the dark matter (DM), is an undeniable fact in terms of observational evidence. Astrophysical and cosmological observational evidence, notably the study of galaxy clusters by Fritz Zwicky Zwicky:1933gu , the rotation curves of galaxy clusters by Vera Rubin Rubin:1970zza and the observation of bullet cluster by Chandra observatory Clowe:2006eq have provided compelling evidence for the existence of this mysterious species in the present universe. Additionally, gravitational lensing and anisotropy of cosmic microwave background radiation (CMBR) measured by several cosmological experiments WMAP:2006bqn ; Roszkowski:2017nbc ; Planck:2018vyg ; Bertone:2004pz ; Hu:2001bc have further supported the need for such enigmatic matter. The DM number density ($\Omega_{DM}h^{2}$) as reported by WMAP WMAP:2006bqn and PLANCK Planck:2018vyg is precisely measured to be $\Omega_{DM}h^{2}=0.12\pm 0.001$ at $68\%$ CL. However, the nature of DM remains a mystery. Though the observational evidence tells us about the gravitational interaction of the DM, the particle origin of it presents an intriguing possibility. If DM is an elementary particle, it is expected to exhibit some level of interaction with the Standard Model (SM). The natural approach would be to consider the weak interaction of DM particles with the SM bath. As an interesting fact, weakly interacting massive particles (WIMPs) of electroweak mass scale can be produced thermally in the early universe and then they can freeze out leaving the thermal relic density close to the DM abundance of the present day. This coincidence is often noted as WIMP miracle Kolb:1990vq .
@@
Other than the WIMP scenario, depending on the interaction strength, DM can be a FIMP (feebly interacting massive particle) Hall:2009bx or a SIMP (Strongly interacting massive particle) PhysRevLett.113.171301 ; PhysRevD.41.2388 in nature. There are various other approaches as well namely super WIMP Feng:2003xh ; Pospelov:2008jk , secluded WIMP Pospelov:2007mp , forbidden DM DAgnolo:2020mpt . However, due to favourable detection possibilities at direct, indirect, and collider search experiments, WIMP is the most popular choice. Direct detection (DD) experiments like LUX LUX:2016ggv , XENONnT XENON:2018voc ; XENON:2023cxc and PANDAX PandaX-II:2016vec ; PandaX-II:2017hlx provides stringent bound on wide ranges of masses and couplings of WIMP. Recent bound from LZ experiment LZ:2022lsv eliminates larger DM parameter space. WIMP scenario has been studied in various types of model set up. Depending on the choice of model, WIMP can be a Dirac or a Majorana fermion, a scalar (pseudo-scalar) or a vector boson. In comparison to fermions and scalars, vector gauge bosons as DM candidates have received a little lesser attention due to the necessity of adding extra gauge extensions. Vector dark matter (VDM) has been studied in Baek:2012se ; Farzan:2012hh ; Baouche:2021wwa ; Nomura:2020zlm ; Babu:2021hef in various gauge extensions. Specifically, VDM in $U(1)$ extension has been analysed in Duch:2015jta ; Amiri:2022cbv whereas our focus is to study such model in conjunction to an extended scalar sector of the SM augmented by a singlet scalar and a $SU(2)_{L}$ triplet scalar. The presence of the triplet in the scalar sector is necessary to facilitate neutrino mass which is discussed in the next paragraph.
@@
Neutrino mass is one of the most compelling reasons to go beyond the Standard Model (BSM). Neutrino oscillation data from various experiments confirmed that neutrinos have tiny mass $\sim{\cal O}(0.1)$ eV 10.1093/ptep/ptaa104 ; Mohapatra:2005wg . The most popular way to generate such tiny neutrino mass at tree level is the Seesaw mechanism. Based on the particle contents of the model, there are three types of Seesaw mechanisms. Type I seesaw Minkowski:1977sc ; Yanagida:1979as ; Gell-Mann:1979vob ; Mohapatra:1979ia extends the SM particle spectrum by three heavy right handed neutrinos where as type II MAGG198061 ; LAZARIDES1981287 ; PhysRevD.23.165 ; PhysRevD.22.2227 ; PhysRevD.22.2860 ; PhysRevLett.44.912 ; Ma:1998dn adds a $SU(2)_{L}$ triplet scalar and type III Foot:1988aq ; Ma:1998dx ; Ma:2002pf ; Hambye:2003rt adds a $SU(2)_{L}$ triplet fermion. Type I seesaw models with an extra $U(1)_{X}$ symmetry have been studied in Ref. Okada:2018tgy ; Okada:2019sbb ; Okada:2020evk in the context of dark matter phenomenology. Interestingly, in addition to the added advantage of vacuum stability, the type II seesaw mechanism can generate small neutrino masses with a $SU(2)_{L}$ triplet scalar with a very small vacuum expectation value (VEV) $v_{t}$ whereas the type I seesaw mechanism requires three right-handed Majorana neutrinos with masses in the range $10^{9}-10^{10}$ GeV. Moreover, if we choose the triplet VEV $v_{t}$ to be in the range $1-3$ GeV( $>10^{-4}$ GeV), we can get a mass range for the doubly charged and singly charged scalar that can be probed at colliders Ashanujjaman:2021txz . Fermionic DM with type II seesaw in $U(1)_{X}$ model has been discussed in Okada:2022cby ; Ghosh:2021khk .
@@
In this article, we explore the possibility of the existence of a gauge boson VDM in a general gauged $U(1)_{X}$ extension of SM where neutrino mass is generated by the type II seesaw mechanism. In addition to a triplet scalar, a $SU(2)_{L}$ singlet, the only particle that carries the $U(1)_{X}$ charge, is necessary to break this symmetry spontaneously resulting in a massive vector boson, $Z^{\prime}$ which plays the role of DM in this model. The stability of DM is ensured by an exact charge conjugation symmetry which arises due to $U(1)_{X}$ invariance. The interaction of the singlet scalar with the SM Higgs doublet acts as a portal between the dark sector and the visible sector. However, the presence of a triplet scalar in this scenario also gives rise to interesting DM phenomenology by mixing with the SM Higgs doublet and the scalar singlet. Therefore we study an interesting scenario that addresses the non-zero neutrino mass and DM abundance in a unified framework. Another interesting factor that prompts us to explore this scenario is the recent combined report by the ATLAS and CMS collaboration on the $h\to Z\gamma$ decay channel. Their findings reveal a combined signal strength $\mu_{Z\gamma}=2.2\pm 0.7$ ATLAS:2023yqk , indicating a slight $2\sigma$ discrepancy compared to the Standard Model prediction of $\mu_{Z\gamma}=1$. Some recent studies Barducci:2023zml ; Boto:2023bpg show a way to accommodate such excesses by different model perspective. Type II seesaw triplet (the singly charged and doubly charged component) can also accommodate non-standard values of $\mu_{Z\gamma}$ as shown in various studies Arbabifar:2012bd ; BhupalDev:2013xol . We show the effect of the triplet charged scalars with the new couplings introduced by the complex scalar singlet in this direction.
@@
This article is organized as follows. In Sec.II, we describe the model in detail along with the generation of neutrino mass by the type II seesaw mechanism. In Sec. III, the experimental and theoretical constraints on the model parameters are discussed. In Sec. IV, we finally discuss the phenomenology of the thermal VDM with the prediction of possible experimental signature of this model parameter space (relic density satisfied) from DD. In Sec. V, the possibility of freeze-in scenario of the VDM is discussed briefly. Finally, the conclusion is drawn in Sec. VI.
@@ II.1 Scalar Sector
$m^{2}_{W}=\frac{1}{4}g^{2}(v^{2}_{d}+2v^{2}_{t}),~{}~{}~{}~{}m^{2}_{Z}=\frac{1}{4}\frac{g^{2}}{cos^{2}\theta_{W}}(v^{2}_{d}+4v^{2}_{t}).$ (10)
where, $\theta_{W}$ is the Weinberg angle. The triplet VEV ($v_{t}$) gives an additional contribution to both $W^{\pm}$ mass and $Z$ boson mass and in turn contributes to $\rho$ parameter (discussed in III). The minimization of the scalar potential yields the following relations:
@@
$m^{2}_{H^{+}}=\frac{v_{d}^{2}+2v_{t}^{2}}{4v_{t}}(2\sqrt{2}\mu-v_{t}\lambda_{4}).$ (18)
@@
In the above equations, $s_{\alpha_{i}}$ and $c_{\alpha_{i}}$ where $i=1,~{}2,~{}3$, stand for $\sin\alpha_{i}$ and $\cos\alpha_{i}$ respectively. Scalar potential parameters ($\lambda$s) can be expressed in terms of the physical masses of the scalars, mixing angles and VEVs. Such relations, listed in Appendix A.1, help us to parameterize the model in terms of physical quantities. So the set of free parameters of this model is $\rm{sin\alpha_{1}}$, $\rm{sin\alpha_{2}}$, $\rm{sin\alpha_{3}}$, $m_{H_{1}}$, $m_{H_{2}}$, $m_{A}$, $m_{H^{\pm\pm}}$, $m_{H^{\pm}}$, $v_{s}$}. We will later see that the triplet scalar masses would be almost degenerate from theoretical constraints. So the set of free parameters are
@@
where $m_{\Delta}$ represents masses of the scalars $m_{H^{++}},m_{H^{+}},m_{H_{1}},m_{A}$ which have large triplet component in them and $m_{S}$ is the mass of scalar $H_{2}$ which is dominantly singlet.
@@ II.2 Dark Sector
The $U(1)_{X}$ invariant Lagrangian enjoys an additional dark conjugation symmetry Farzan:2012hh ; Ma:2017ucp described as
@@ II.3 Yukawa Lagrangian and neutrino mass
where $Y_{\Delta}$ is the $3\times 3$ Yukawa matrix, $C$ is the charge conjugation matrix and $L$ are the SM lepton doublet. The following Majorana mass for neutrinos can be generated from the above Yukawa term,
@@
Since $v_{t}\sim\mathcal{O}(GeV)$ in our analysis and the neutrino mass scale is in the eV range, therefore $Y_{\Delta}$ must be very small. For neutrino mass scale $m_{\nu}\sim 0.8$ eV and $v_{t}\sim 1$ GeV, $Y_{\Delta}\sim 5.66\times 10^{-10}$. Apart from the kinematics, the value of $Y_{\Delta}$ controls the decay mode of $H^{\pm\pm}$. The $\Gamma(H^{\pm\pm}\to\ell^{\pm}\ell^{\pm})\propto{\mid Y_{\Delta}\mid}^{2}$, thus for our choice of $v_{t}$ which corresponds to very small $Y_{\Delta}$, the like signed dilepton decay of $H^{\pm\pm}$ would be highly suppressed and $W^{\pm}W^{\pm}$ channel will dominate.
@@ III Theoretical and experimental constraints on model parameters
Now we are ready to discuss possible experimental and theoretical constraints on the parameters of this model which would be used in our analysis.
@@ III.1 Experimental constraints
$\bullet$ $\rho$ parameter constraint: The triplet vacuum expectation value $v_{t}$ contributes to the $W^{\pm}$ and $Z$ boson masses at the tree-level (see Eqn. 10), thus affecting the value of the $\rho$ parameter which is defined in terms of the s $v_{d}$ and $v_{t}$ of the SM doublet and triplet scalars respectively as
@@
$\bullet$ Experimental limits on singly and doubly charged scalar masses: This scenario predicts a plethora of heavy scalars that have been extensively searched at both $e^{+}e^{-}$ and hadron colliders. The negative search results at LEP-2 puts a limit on the singly charged mass $m_{H^{\pm}}>78$ (GeV) LEPHiggsWorkingGroupforHiggsbosonsearches:2001ogs . On the other hand, the lower limit on the doubly charged scalar mass $m_{H^{\pm\pm}}$ from the LHC is highly model dependent. The decay modes of the $H^{\pm\pm}$ depend on $v_{t}$ and the mass splitting $m_{H^{\pm\pm}}-m_{H^{\pm}}$. For $v_{t}<10^{-4}$ GeV (large Yukawa coupling) and assuming degenerate scalars, $H^{\pm\pm}$ decays to like-signed dilepton with almost $100\%$ probability. The decay mode of $H^{\pm\pm}$ changes rather drastically when $v_{t}>10^{-4}$ GeV, leading to various competing channels, like $H^{\pm}H^{\mp}$, $W^{\pm}W^{\mp}$, and $W^{\pm}H^{\mp}$, if kinematically allowed. From current analysis at the LHC, $m_{H^{\pm\pm}}\gtrsim 1080$ GeV ATLAS:2022pbd is allowed for $H^{\pm\pm}\to\ell^{\pm}\ell^{\pm}$ decay where the branching ratios to each of the possible leptonic final states ($ee,\mu\mu,\tau\tau,e\mu,e\tau,\mu\tau$) are considered to be equal. However in a type II seesaw framework, when $H^{\pm\pm}\to W^{\pm}W^{\pm}$ decay is allowed, doubly charged Higgs mass has a lower bound of 420 GeV Ashanujjaman:2021txz . In the following analysis, we have chosen $v_{t}\sim\mathcal{O}$(1 GeV) and $m_{H^{\pm\pm}}$ to be greater than 420 GeV.
$\bullet$ Collider constraints on the neutral scalar masses and mixing: The Higgs decay width measurement at LHC can constrain the mixing angle between the triplet and SM doublet. The upper bound on this mixing angle, $\sin\alpha_{1}$ is $\lesssim 0.05$ Bhattacharya:2017sml which is consistent with the experimental observation of $h\to WW^{*}$ branching ratio. A similar bound is obtained on the mixing angle of the complex scalar singlet with SM Higgs from theoretical and experimental constraints. The most stringent bound comes from W mass correction at NLO Lopez-Val:2014jva . For $m_{S}$ in the range $\sim\{200-1000\}$ GeV, the upper bound on the mixing angle $\sin\alpha_{2}$ varies between ($0.2-0.3$) Robens:2016xkb . In this model, the value of $\sin\alpha_{1}$ is taken to be $<0.05$ and the maximum value of $\sin\alpha_{2}$ is taken to be $0.1$.
@@
Among all possible combinations from the above-mentioned branching ratios, the most stringent bound comes from $\mu\to eee$ with BR($\mu\to eee$) $\lesssim 1.0\times 10^{-12}$ from SINDRUM experiment SINDRUM:1987nra . As Yukawa parameters are inversely proportional to $v_{t}$, we can extract a lower limit on $v_{t}$ from the above expression for a particular mass of $H^{++}$. Another lepton flavor violating process, namely $\ell_{i}\to\ell_{j}\gamma$ receives non-standard contribution via loop diagrams of charged (singly and doubly) scalars. The BR of $\mu\to e\gamma$ gives the most stringent bound among these decays and the corresponding expression is given by Okada:2022cby
@@
The latest bound on ${\rm BR}(\mu\to e\gamma)$ ($\lesssim 4.2\times 10^{-13}$) has been recently reported by MEG collaboration MEG:2016leq . In this model, $Y_{\Delta}$ being very small as mentioned before (see under Eqn. 30), the branching ratios of the aforementioned processes are way smaller than the reported experimental bounds.
$\bullet$ Constraint from oblique parameters: Dominant contribution in oblique parameters comes from the charged Higgs sector of the triplet scalar. The most stringent bound comes from $T$ parameter and it demands the mass splitting of the doubly charged and the singly charged scalar to be $\lesssim$ 40 GeV Chun:2012jw . Our DM sector constraints are insensitive to this mass splitting, so the triplet scalar masses are taken to be degenerate for the analysis of the DM part.
$\bullet$ Higgs Invisible decay constraint: Current Higgs searches at the LHC puts stringent limit on the branching fraction of Higgs invisible decay. In our scenario, SM Higgs can decay to a pair of VDM through mixing with the singlet scalar when $m_{h}>2M_{Z^{\prime}}$. The expression of Higgs invisible branching ratio is given by

### v2

@@ Abstract
We study the phenomenology of a vector dark matter (VDM) in a $U(1)_{X}$ gauged extension of the Standard Model (SM) which is set in a type II seesaw framework. When this $U(1)_{X}$ symmetry is spontaneously broken by the vacuum expectation value (VEV) of a newly introduced complex scalar singlet, the gauge boson $Z^{\prime}$ becomes massive. The stability of the dark matter (DM) is ensured by the presence of an exact charge conjugation symmetry resulting from the structure of the Lagrangian. On the other hand, the $SU(2)_{L}$ triplet scalar facilitates light neutrino masses via the type II seesaw mechanism. We have studied the phenomenology of the usual WIMP DM, considering all possible theoretical and experimental constraints that are applicable. Due to the presence of a triplet scalar, the present framework can also accommodate the observed $2\sigma$ deviation in $h\to Z\gamma$ decay rate, recently measured at the LHC. The possibility of non-thermal production of DM from the decay of the singlet scalar has also been briefly discussed.
@@ I Introduction
The existence of a non-luminous, non-baryonic form of matter which provides roughly $26\%$ energy density of the universe, popularly known as the dark matter (DM), is an undeniable fact in terms of observational evidence. Astrophysical and cosmological observational evidence, notably the study of galaxy clusters by Fritz Zwicky Zwicky (1933), the rotation curves of galaxy clusters by Vera Rubin Rubin and Ford (1970) and the observation of bullet cluster by Chandra observatory Clowe et al. (2006) have provided compelling evidence for the existence of this mysterious species in the present universe. Additionally, gravitational lensing and anisotropy of cosmic microwave background radiation (CMBR) measured by several cosmological experiments Spergel et al. (2007); Roszkowski et al. (2018); Aghanim et al. (2020); Bertone et al. (2005); Hu and Dodelson (2002) have further supported the need for such enigmatic matter. The DM number density ($\Omega_{DM}h^{2}$) as reported by WMAP Spergel et al. (2007) and PLANCK Aghanim et al. (2020) is precisely measured to be $\Omega_{DM}h^{2}=0.12\pm 0.001$ at $68\%$ CL. However, the nature of DM remains a mystery. Though the observational evidence tells us about the gravitational interaction of the DM, the particle origin of it presents an intriguing possibility. If DM is an elementary particle, it is expected to exhibit some level of interaction with the Standard Model (SM). The natural approach would be to consider the weak interaction of DM particles with the SM bath. As an interesting fact, weakly interacting massive particles (WIMPs) of electroweak mass scale can be produced thermally in the early universe and then they can freeze out leaving the thermal relic density close to the DM abundance of the present day. This coincidence is often noted as WIMP miracle Kolb and Turner (1990).
@@
Other than the WIMP scenario, depending on the interaction strength, DM can be a FIMP (feebly interacting massive particle) Hall et al. (2010) or a SIMP (Strongly interacting massive particle) Hochberg et al. (2014); Dimopoulos et al. (1990) in nature. There are various other approaches as well namely super WIMP Feng et al. (2003); Pospelov et al. (2008a), secluded WIMP Pospelov et al. (2008b), forbidden DM D’Agnolo et al. (2021). However, due to favourable detection possibilities at direct, indirect, and collider search experiments, WIMP is the most popular choice. Direct detection (DD) experiments like LUX Akerib et al. (2017), XENONnT Aprile et al. (2018, 2023) and PANDAX Tan et al. (2016); Cui et al. (2017) provides stringent bound on wide ranges of masses and couplings of WIMP. Recent bound from LZ experiment Aalbers et al. (2023) eliminates larger DM parameter space. WIMP scenario has been studied in various types of model set up. Depending on the choice of model, WIMP can be a Dirac or a Majorana fermion, a scalar (pseudo-scalar) or a vector boson. In comparison to fermions and scalars, vector gauge bosons as DM candidates have received a little lesser attention due to the necessity of adding extra gauge extensions. Vector dark matter (VDM) has been studied in Baek et al. (2013); Farzan and Akbarieh (2012); Baouche et al. (2021); Nomura et al. (2021); Babu et al. (2022); Barman et al. (2017) in various gauge extensions. Specifically, VDM in $U(1)$ extension has been analysed in Duch et al. (2015); Amiri et al. (2022). In the above scenario, the dark matter couples to the SM through the Higgs portal via the mixing between the SM Higgs and the new scalar singlet. Here, we have discussed the variation of the VDM parameter space where the scalar sector has been extended with an additional triplet scalar. $SU(2)_{L}$ triplet Higgs seems to be a very obvious candidate in this context as the presence of a triplet in the scalar sector can facilitate neutrino mass thereby solving another important question of particle physics. The neutrino mass generation will be discussed in the next paragraph. The presence of the extra triplet will provide extra annihilation channels for the dark matter, thus it has more control in tuning of the relic abundance.
@@
Neutrino mass is one of the most compelling reasons to go beyond the Standard Model (BSM). Neutrino oscillation data from various experiments confirmed that neutrinos have tiny mass $\sim{\cal O}(0.1)$ eV Group et al. (2020); Mohapatra et al. (2007). The most popular way to generate such tiny neutrino mass at tree level is the Seesaw mechanism. Based on the particle contents of the model, there are three types of Seesaw mechanisms. Type I seesaw Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979); Mohapatra and Senjanovic (1980) extends the SM particle spectrum by three heavy right handed neutrinos where as type II Magg and Wetterich (1980); Lazarides et al. (1981); Mohapatra and Senjanović (1981); Schechter and Valle (1980); Cheng and Li (1980); Mohapatra and Senjanović (1980); Ma (1998) adds a $SU(2)_{L}$ triplet scalar and type III Foot et al. (1989); Ma and Sarkar (1998); Ma and Roy (2002); Hambye et al. (2004) adds a $SU(2)_{L}$ triplet fermion. Type I seesaw models with an extra $U(1)_{X}$ symmetry have been studied in Ref. Okada et al. (2019); Okada and Seto (2020); Okada et al. (2020) in the context of dark matter phenomenology. Interestingly, in addition to the added advantage of vacuum stability, the type II seesaw mechanism can generate small neutrino masses with a $SU(2)_{L}$ triplet scalar with a very small vacuum expectation value (VEV) $v_{t}$ whereas the type I seesaw mechanism requires three right-handed Majorana neutrinos with masses in the range $10^{9}-10^{10}$ GeV. Moreover, if we choose the triplet VEV $v_{t}$ to be in the range $1-3$ GeV( $>10^{-4}$ GeV), we can get a mass range for the doubly charged and singly charged scalar that can be probed at colliders Ashanujjaman and Ghosh (2022). Fermionic DM with type II seesaw in $U(1)_{X}$ model has been discussed in Okada and Seto (2022); Ghosh et al. (2022).
@@
In this article, we explore the possibility of the existence of a gauge boson VDM in a general gauged $U(1)_{X}$ extension of SM where neutrino mass is generated by the type II seesaw mechanism. In addition to a triplet scalar, a $SU(2)_{L}$ singlet, the only particle that carries the $U(1)_{X}$ charge, is necessary to break this symmetry spontaneously resulting in a massive vector boson, $Z^{\prime}$ which plays the role of DM in this model. The stability of DM is ensured by an exact charge conjugation symmetry which arises due to $U(1)_{X}$ invariance. The interaction of the singlet scalar with the SM Higgs doublet acts as a portal between the dark sector and the visible sector. Moreover, the presence of a triplet scalar in this scenario gives rise to interesting DM phenomenology by mixing with the SM Higgs doublet and the scalar singlet and finally resulting into an enhanced DM parameter space. Therefore we study an interesting scenario that addresses the non-zero neutrino mass and DM abundance in a unified framework. Another interesting factor that prompts us to explore this scenario is the recent combined report by the ATLAS and CMS collaboration on the $h\to Z\gamma$ decay channel. Their findings reveal a combined signal strength $\mu_{Z\gamma}=2.2\pm 0.7$ Aad et al. (2024), indicating a slight $2\sigma$ discrepancy compared to the Standard Model prediction of $\mu_{Z\gamma}=1$. Some recent studies Barducci et al. (2023); Boto et al. (2023) show a way to accommodate such excess by different model perspectives. Type II seesaw triplet (the singly charged and doubly charged component) can also accommodate non-standard values of $\mu_{Z\gamma}$ as shown in various studies Arbabifar et al. (2013); Bhupal Dev et al. (2013). We show that in addition to the dominant contribution of the triplet charged scalar components, the new couplings introduced by the complex scalar singlet can affect the signal strengths $\mu_{\gamma\gamma}$ and $\mu_{Z\gamma}$.
@@
This article is organized as follows. In Sec.II, we describe the model in detail along with the generation of neutrino mass by the type II seesaw mechanism. In Sec. III, the experimental and theoretical constraints on the model parameters are discussed. In Sec. IV, we finally discuss the phenomenology of the thermal VDM with the prediction of possible experimental signature of this allowed model parameter space from DD. In Sec. V, the possibility of freeze-in scenario of the VDM is discussed briefly. Finally, the conclusion is drawn in Sec. VI.
@@ II.1 Scalar Sector
$m^{2}_{W}=\frac{1}{4}g^{2}(v^{2}_{d}+2v^{2}_{t}),~{}~{}~{}~{}m^{2}_{Z}=\frac{1}{4}\frac{g^{2}}{cos^{2}\theta_{W}}(v^{2}_{d}+4v^{2}_{t}),~{}~{}M^{2}_{Z^{\prime}}=g^{2}_{x}v^{2}_{s}.$ (10)
where, $\theta_{W}$ is the Weinberg angle. The triplet VEV ($v_{t}$) gives an additional contribution to both $W^{\pm}$ mass and $Z$ boson mass and in turn contributes to $\rho$ parameter (discussed in Sec. III). The minimization of the scalar potential yields the following relations:
@@
$m^{2}_{H^{\pm}}=\frac{v_{d}^{2}+2v_{t}^{2}}{4v_{t}}(2\sqrt{2}\mu-v_{t}\lambda_{4}).$ (18)
@@
In the above equations, $s_{\alpha_{i}}$ and $c_{\alpha_{i}}$ where $i=1,~{}2,~{}3$, stand for $\sin\alpha_{i}$ and $\cos\alpha_{i}$ respectively. Scalar potential parameters ($\lambda$s) can be expressed in terms of the physical masses of the scalars, mixing angles, and VEVs. Such relations, listed in Appendix A.1, help us to parameterize the model in terms of physical quantities. So the set of free parameters of this model is $\rm{sin\alpha_{1}}$, $\rm{sin\alpha_{2}}$, $\rm{sin\alpha_{3}}$, $m_{H_{1}}$, $m_{H_{2}}$, $m_{A}$, $m_{H^{\pm\pm}}$, $m_{H^{\pm}}$, $v_{s}$}. We will later see that the triplet scalar masses would be almost degenerate from theoretical constraints. So the set of free parameters are
@@
where $m_{\Delta}$ represents masses of the scalars $m_{H^{\pm\pm}},m_{H^{\pm}},m_{H_{1}},m_{A}$ which have large triplet component in them and $m_{S}$ is the mass of scalar $H_{2}$ which is dominantly singlet.
@@ II.2 Dark Sector
The $U(1)_{X}$ invariant Lagrangian enjoys an additional dark conjugation symmetry Farzan and Akbarieh (2012); Ma (2017) described as
@@ II.3 Yukawa Lagrangian and neutrino mass
where $Y_{\Delta}$ is the $3\times 3$ Yukawa matrix, $C$ is the charge conjugation matrix and $L$ is the SM lepton doublet. The following Majorana mass for neutrinos can be generated from the above Yukawa term,
@@
Since $v_{t}\sim\mathcal{O}(GeV)$ in our analysis and the neutrino mass scale is in the eV range, therefore $Y_{\Delta}$ must be very small. For neutrino mass scale $m_{\nu}\sim 0.8$ eV and $v_{t}\sim 1$ GeV, $Y_{\Delta}\sim 5.66\times 10^{-10}$. Apart from the kinematics, the value of $Y_{\Delta}$ controls the decay mode of $H^{\pm\pm}$. The partial decay width of $H^{\pm\pm}$ to like sign dileptons being proportional to ${\mid Y_{\Delta}\mid}^{2}$, $H^{\pm\pm}\to\ell^{\pm}\ell^{\pm}$ would be highly suppressed for our choice of $v_{t}$ which corresponds to very small $Y_{\Delta}$ and $H^{\pm\pm}\to W^{\pm}W^{\pm}$ channel will dominate consequently.
@@ III Theoretical and experimental constraints on model parameters
Now we are ready to discuss the possible experimental and theoretical constraints on the parameters of this model which would be used in our analysis.
@@ III.1 Experimental constraints
$\bullet$ $\rho$ parameter constraint: The triplet vacuum expectation value $v_{t}$ contributes to the $W^{\pm}$ and $Z$ boson masses at the tree-level (see Eqn. 10), thus affecting the value of the $\rho$ parameter which is defined in terms of the VEVs $v_{d}$ and $v_{t}$ of the SM doublet and triplet scalars respectively as
@@
$\bullet$ Experimental limits on singly and doubly charged scalar masses: This scenario predicts a plethora of heavy scalars that have been extensively searched at both $e^{+}e^{-}$ and hadron colliders. The negative search results at LEP-2 puts a limit on the singly charged mass $m_{H^{\pm}}>78$ (GeV) LEP (2001). On the other hand, the lower limit on the doubly charged scalar mass $m_{H^{\pm\pm}}$ from the LHC is highly model dependent. The decay modes of the $H^{\pm\pm}$ depend on $v_{t}$ and the mass splitting $m_{H^{\pm\pm}}-m_{H^{\pm}}$. For $v_{t}<10^{-4}$ GeV (large Yukawa coupling) and assuming degenerate scalars, $H^{\pm\pm}$ decays to like-signed dilepton with almost $100\%$ probability. The decay mode of $H^{\pm\pm}$ changes rather drastically when $v_{t}>10^{-4}$ GeV, leading to various competing channels, like $H^{\pm}H^{\mp}$, $W^{\pm}W^{\mp}$, and $W^{\pm}H^{\mp}$, if kinematically allowed. From current analysis at the LHC, $m_{H^{\pm\pm}}\gtrsim 1080$ GeV Aad et al. (2023a) is allowed for $H^{\pm\pm}\to\ell^{\pm}\ell^{\pm}$ decay where the branching ratios to each of the possible leptonic final states ($ee,\mu\mu,\tau\tau,e\mu,e\tau,\mu\tau$) are considered to be equal. However in a type II seesaw framework, when $H^{\pm\pm}\to W^{\pm}W^{\pm}$ decay is allowed, doubly charged Higgs mass has a lower bound of 420 GeV Ashanujjaman and Ghosh (2022). In the following analysis, we have chosen $v_{t}\sim\mathcal{O}$(1 GeV) and $m_{H^{\pm\pm}}$ to be greater than 420 GeV.
$\bullet$ Collider constraints on the neutral scalar masses and mixing: The Higgs decay width measurement at LHC can constrain the mixing angle between the triplet and SM doublet. The upper bound on this mixing angle, $\sin\alpha_{1}$ is $\lesssim 0.05$ Bhattacharya et al. (2017) which is consistent with the experimental observation of $h\to WW^{*}$ branching ratio. A similar bound is obtained on the mixing angle of the complex scalar singlet with SM Higgs from theoretical and experimental constraints. The most stringent bound comes from W mass correction at NLO López-Val and Robens (2014). For $m_{S}$ in the range $\sim\{200-1000\}$ GeV, the upper bound on the mixing angle $\sin\alpha_{2}$ varies between ($0.2-0.3$) Robens and Stefaniak (2016). In this model, the value of $\sin\alpha_{1}$ is taken to be $<0.05$ and the maximum value of $\sin\alpha_{2}$ is taken to be $0.1$.
@@
Among all possible combinations from the above-mentioned branching ratios, the most stringent bound comes from $\mu\to eee$ with BR($\mu\to eee$) $\lesssim 1.0\times 10^{-12}$ from SINDRUM experiment Bellgardt et al. (1988). As Yukawa parameters are inversely proportional to $v_{t}$, we can extract a lower limit on $v_{t}$ from the above expression for a particular mass of $H^{\pm\pm}$. Another lepton flavor violating process, namely $\ell_{i}\to\ell_{j}\gamma$ receives non-standard contribution via loop diagrams of charged (singly and doubly) scalars. The BR of $\mu\to e\gamma$ gives the most stringent bound among these decays and the corresponding expression is given by Okada and Seto (2022)
@@
The latest bound on ${\rm BR}(\mu\to e\gamma)$ ($\lesssim 4.2\times 10^{-13}$) has been recently reported by MEG collaboration Baldini et al. (2016). In this model, $Y_{\Delta}$ being very small as mentioned before (see under Eqn. 30), the branching ratios of the aforementioned processes are way smaller than the reported experimental bounds.
$\bullet$ Constraint from oblique parameters: Dominant contribution in oblique parameters comes from the charged Higgs sector of the triplet scalar. The most stringent bound comes from the $T$ parameter and it demands the mass splitting of the doubly charged and the singly charged scalars to be $\lesssim$ 40 GeV Chun et al. (2012). Our DM sector constraints are insensitive to this mass splitting, so the triplet scalar masses are taken to be degenerate for the analysis of the DM part.
$\bullet$ Higgs Invisible decay constraint: Current Higgs searches at the LHC puts stringent limits on the branching fraction of Higgs invisible decay. In our scenario, SM Higgs can decay to a pair of VDM through mixing with the singlet scalar when $m_{h}>2M_{Z^{\prime}}$. The expression of Higgs invisible branching ratio is given by